OpenStudy (anonymous):

I really need help!! which point represents a solution to the system y≤1/5x y>4x-3 Answers: (5,1) (0,0) (-2,-3) (2,5)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it is usually C want to check that one first ?

OpenStudy (misty1212):

oh maybe it is A lets check that first

OpenStudy (anonymous):

I checked but Im not sure because it is supposed to be 2 answers

OpenStudy (misty1212):

y≤1/5x for \(x=5,y=1\) get \(1\leq \frac{5}{5}\) which is true

OpenStudy (misty1212):

and \[1>4\times 5-1\]which is false, so not that one

OpenStudy (misty1212):

looks like \((0,0)\) works for both

OpenStudy (misty1212):

\[-3\leq \frac{-2}{5}\] is true

OpenStudy (misty1212):

\[-3>4\times -2-3\] is also true

OpenStudy (misty1212):

so looks like B and C